Calculate the electric flux for each of the following sets of variables.

a. Electric Field = 7.06 x 1013N/C

A = 4.4 x 10-6 m2

θ = 0°

b. Electric Field = 6.9 x 103 N/C

A = 3.1 x 10-3 m2

θ = 45°

c. Electric Field = 8.2 x 105 N/C

A = 6.1 x 10-7 m2

θ = 90°


1
Expert's answer
2022-03-08T08:56:51-0500

The electric flux through the surface is given by

"\Phi=EA\cos\theta"

(a)

"\Phi=7.06*10^{13}* 4.4 *10^{-6}\cos0^\circ\\
=3.11*10^{8}\:\rm N\cdot m^2/C"

(b)

"\Phi=6.9*10^{3}* 3.1 *10^{-3}\cos45^\circ\\
=15\:\rm N\cdot m^2/C"

(c)

"\Phi=8.2*10^{5}* 6.1 *10^{-7}\cos90^\circ\\
=0\:\rm N\cdot m^2/C"